OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out first a red, then a blue marble if none are replaced 7/132 1/11 12/23 7/23

OpenStudy (lucaz):

you have a total of 12 marbles, then the probability P(red) = 3/12

OpenStudy (lucaz):

now the total is 12-1red = 11, then the P(b) = 4/11

OpenStudy (lucaz):

multiply these probabilities 3/12 * 4/11

OpenStudy (lucaz):

@kbear6812

OpenStudy (anonymous):

OKAY IM CALCULATING

OpenStudy (anonymous):

i got 1/11

OpenStudy (lucaz):

you got it right